What are the mole fractions of solute and solvent in a 1.00 molal solution
1
Expert's answer
2018-10-08T07:52:09-0400

Solution:

1.n(H2O)=m/M;m=1000 gram;M=18 gram/mole;

n(H2O)=1000/18=55.556 mol;

2.Mole fractions of solute: 1/(1+55.556)=0.018

3.Mole fractions of solvent: 1-0.117=0.982

Answer: 0.018;0.982.